OpenStudy (anonymous):

Medal! How can this multiplication problem be written using only improper fractions? 5 x 2 1/8 = ?

OpenStudy (lacypennelll):

Multiple choice?

OpenStudy (anonymous):

Yes

OpenStudy (lacypennelll):

5/1 x 17/8

OpenStudy (nincompoop):

\[a \frac{ b }{ c } = c \times a + b\]

OpenStudy (anonymous):

Explain it to me...?

OpenStudy (nincompoop):

well try to think about it first and attempt to understand what it may be

OpenStudy (anonymous):

Ok thank you, I'll see if I can get it right brb.

OpenStudy (anonymous):

Wow I seriously didn't read that. I just had to simplify the improper fraction. 8 x 2 + 1.

OpenStudy (anonymous):

Thank you for help!

OpenStudy (nincompoop):

\[10 \frac{ 2 }{ 3 } \rightarrow a=10, b=2, c=3 \] \[c \times a + b = 3 \times 10 + 2\]

OpenStudy (anonymous):

Thank you for the medal. <3

OpenStudy (nincompoop):

you're welcome